# Would You Trust an AI Doctor? Building Reliable Medical Predictions with Kernel Dropout Uncertainty

### v1

@@ Abstract
The growing capabilities of AI raise questions about their trustworthiness in healthcare, particularly due to opaque decision-making and limited data availability. This paper proposes a novel approach to address these challenges, introducing a Bayesian Monte Carlo Dropout model with kernel modelling. Our model is designed to enhance reliability on small medical datasets, a crucial barrier to the wider adoption of AI in healthcare. This model leverages existing language models for improved effectiveness and seamlessly integrates with current workflows. We demonstrate significant improvements in reliability, even with limited data, offering a promising step towards building trust in AI-driven medical predictions and unlocking its potential to improve patient care. Our source codes are freely and publicly available at:
Would You Trust an AI Doctor? Building Reliable Medical Predictions with Kernel Dropout Uncertainty
Ubaid Azam1, Imran Razzak2, Shelly Vishwakarma1, Hakim Hacid3, Dell Zhang4, and Shoaib Jameel1 1 School of electronics and Computer Science, University of Southampton, United Kindom 2School of Computer Science and Engineering at University of New South Wales, Sydney, Australia 3Technology Innovation Institute, UAE, 4Thomson Reuters Labs, UK {u.azam,s.vishwakarma,M.S.Jameel}@soton.ac.uk imran.razzak@unsw.edu.au hakim.hacid@tii.ae dell.z@ieee.org
@@ 1 Introduction
The current healthcare landscape is witnessing a remarkable transformation driven by machine learning Rogan et al. (2024); Chen et al. (2024). Automated methods are proliferating across diverse medical domains, empowering tasks like medical imaging analysis with the ability to detect anomalies Bercea et al. (2024), classify diseases like cancer Chanda et al. (2024a), and even predict disease progression Young et al. (2024). However, despite the undeniable power of these tools, responsible implementation necessitates careful consideration of ethical issues Weidener et al. (2024), data privacy Zhou et al. (2024), and the crucial role of human oversight. One persistent challenge lies in ensuring the consistently high reliability Jerez et al. (2022); Singpurwalla (2006) of automated predictions, a factor upon which the trust and adoption of these models by medical practitioners ultimately hinge Chanda et al. (2024b); Joshi et al. (2024); Wani et al. (2024).
@@
Another challenge is developing computational methods that are reliable even on small datasets. In the medical domain, datasets are typically scarce, even those publicly available Spasic et al. (2020). This scarcity stems from the limited release of critical private data, a trend driven by strong patient privacy concerns and regulations safeguarding that privacy in many countries Moore and Frye (2020). Restricted access to data often forces training models on smaller datasets, making complete trust in their predictions more challenging. This underscores the critical importance of ensuring the reliability and dependability of these predictions. This paper addresses this challenge by presenting a novel computational model built on Bayesian Monte Carlo Dropout, with the specific objective of enhancing the reliability and trustworthiness of predictions made in the medical domain.
@@
Despite advancements, machine learning approaches in healthcare face challenges in building trust due to limitations in their transparency and reliability Damianou and Lawrence (2013). For example, deep learning, while powerful, often functions as a “black box”, making it difficult to understand its reasoning and assess its accuracy in individual cases. In contrast, deep Bayesian models Bui et al. (2016); Li et al. (2021); Abdullah et al. (2022); Abdar et al. (2021) provide a wide range of benefits Peng et al. (2019a); Li et al. (2021), for instance, in the medical predictions it could boast of several potential advantages Yang and Shafto (2017); Guo et al. (2018); Dolezal et al. (2022). A Bayesian model inherently provides uncertainty estimates with predictions, giving clinicians a clearer understanding of how confident the model is in its output. This allows for better-informed decision-making while acknowledging the limitations of the model. By averaging over multiple model configurations (due to dropout), the Bayesian model can be less susceptible to noise and biases present in training data, leading to more generalizable and reliable predictions. The Bayesian setting outputs represent probability distributions, allowing clinicians to understand the range of possible outcomes and the likelihood of each. This aids in explaining the model’s reasoning and building trust in its predictions. Analyzing dropout patterns can offer clues about which features influence the model’s predictions most, potentially revealing previously unknown relationships within the data. Bayesian modelling can be calibrated to ensure their predictions align with actual outcomes, further improving their trustworthiness.
@@
Studies have indicated that machine learning contributes to enhancing the efficiency of medical-related tasks Rafi et al. (2021). However, there appears to be resistance among people when it comes to adopting medical artificial intelligence treatments Longoni et al. (2019). Similarly, doctors also feel hesitant to fully embrace it due to the sensitive nature of the tasks and the opaque and black box nature of machine learning models  [^1] https://www.ntu.edu.sg/medicine/news-events/dean’s-blog/trust-can-we-trust-the-machine-can-we-trust-the-doctor Hallowell et al. (2022). From a larger perspective, ensuring reliable decisions is crucial, especially when someone’s life is on the line, allowing little room for mistakes. However, machine learning models face a challenge as they cannot measure the certainty of their decisions.
@@
Our key contributions are the following: 1) We develop a novel computational model based on Bayesian deep learning that exploits a novel Bayesian Monte Carlo Dropout mechanism to model uncertainty. 2) Our fully Bayesian model incorporates conjugate priors that allow us to naturally incorporate prior knowledge or beliefs about the parameters we are estimating. This can lead to more accurate and reliable posterior estimates, especially when data is limited. 3) We develop a kernel model to reliably model the features. Kernels can allow us to model the nature of the data and problem, allowing for better flexibility and adaptation. 4) We demonstrate that the novel model is reliable even on small datasets in binary and multi-class settings.
@@ 2 Related Work
The adoption of machine learning in the medical field is becoming increasingly widespread Callahan and Shah (2017); Wiens and Shenoy (2018), particularly with the development of transformer models tailored for medical applications Peng et al. (2019b); Lee et al. (2020). These models have notably improved the effectiveness and productivity of medical tasks Wang et al. (2023a); Nerella et al. (2023). Especially there is rapid research going on in clinical judgment or diagnosis tasks in which the judgement is made based on clinical notes, medical transcriptions and the data based on the patient history Feng et al. (2020); Gao et al. (2023).
@@
In Zhu et al. (2021), the authors provided an overview of the current transformer models in the medical domain and their utilization in a range of medical applications, such as text mining, question answering, and classification. Similarly, Li et al. (2022) highlighted the challenges encountered in managing Electronic Health Records (EHRs) and how advanced natural language processing techniques are aiding in addressing these challenges. These techniques assist in organizing and generating the data, extracting valuable information, and facilitating classification and prediction tasks, thereby enhancing the effectiveness and efficiency of work processes. Considerable research is currently underway in this area, with Kong et al. (2022) notably introducing a novel medical report generation model called TranSQ. This model generates a series of semantic features to align with plausible clinical concerns and constructs the report using sentence retrieval and selection techniques. Their model surpassed benchmark models in performance on the generation tasks. Additionally, Sezgin et al. (2022) investigated the potential of the widely acclaimed Generative Pretrained Transformer 3 (GPT-3) and outlined considerations for its integration into clinical practices.
@@
Similar to other medical tasks, the classification task is also reaping the benefits of advanced machine learning techniques, aiming to improve healthcare and enhance patient care, given the importance of swiftly analyzing text for clinical decision support, research, and process optimization. Yao et al. (2019) employed deep learning models to classify clinical records into five primary disease categories in traditional Chinese medicine. Amin et al. (2019) performed multi-label classification of ICD-10 code problems by using transfer learning in conjunction with pre-trained BERT and BioBERT models. Meanwhile, Singh et al. (2020) utilized a BERT model to identify diagnoses from unstructured clinical text. Similarly, Yogarajan et al. (2021) conducted multi-label classification of medical texts using advanced transformer models. Through their experiments, they concluded that domain-specific transformers yield superior results compared to standard transformers.
@@
There are certain constraints associated with employing these advanced transformer models for clinical text classification, as noted by Gao et al. (2021). One significant limitation is the black-box nature of these deep learning models, which poses challenges in trusting and implementing these models, particularly in healthcare, where decisions are matters of life and death Petch et al. (2022). These opaque models fail to provide a quantifiable measure of their output confidence. Research is underway to address this issue, with Gal and Ghahramani (2016) employing dropout as a Bayesian approximation to represent uncertainty in deep learning models. Tanneru et al. (2023) introduced two novel metrics, namely Verbalized Uncertainty and Probing Uncertainty, to measure the uncertainty of generated explanations by large language models.
@@ 3 Our Novel Model
This section dives into the technical details of our innovative model, which delivers a more reliable quantification of uncertainty in its predictions for multi-class classification. This uncertainty arises from the inherent lack of complete knowledge about the true relationship between input and output data. We address this uncertainty by differentiating our model from the standard Monte Carlo dropout model Gal and Ghahramani (2016); Magris and Iosifidis (2023) in two key ways.
@@
Our second innovation lies in integrating priors within the Monte Carlo dropout framework of Shelmanov et al. (2021). Priors inject vital flexibility by allowing us to leverage our domain knowledge and beliefs about the problem at hand. This strategic infusion of information guides the model towards solutions that resonate with real-world expectations, effectively biasing it towards sensible outcomes. Additionally, incorporating priors empowers our Bayesian models to quantify uncertainty in their predictions. This explicit estimation of uncertainty is crucial for tasks where understanding the model’s limitations is paramount, such as in medical diagnosis or legal prediction. Finally, priors facilitate robust comparison and selection of different models based on their posterior probabilities. This data-driven approach empowers us to identify the model that most effectively aligns with both the observed data and our prior beliefs.
@@
Imagine a neural network constantly questioning its abilities. This is the essence of Monte Carlo dropout, a process where the network throws “curveballs” at itself during training. At each training step, the network randomly deactivates some neurons in its hidden layers, simulating “failures” and forcing others to take over. This introduces variability and prevents over-reliance on specific neurons. By “silencing” neurons, the network effectively asks, “Can I still work even without these?” This builds a more robust network, adaptable to new data and challenges. During prediction, the network performs multiple passes with different silenced neurons, mimicking multiple “runs” with potential errors. This captures the inherent uncertainty in real-world data and allows the network to express its level of confidence in each prediction. This connection between dropout and uncertainty estimation was not always clear. Earlier works like Gal and Ghahramani (2016) focused on dropout as a regularization technique, but later studies (e.g., Khan et al. (2019); MacKay et al. (1998)) revealed its deeper connection to Bayesian inference in Gaussian Processes. By leveraging dropout, we can not only improve a network’s performance but also gain valuable insights into its limitations and confidence in its predictions.
@@
$q(\mathbf{c}^{*}|\mathbf{x}^{*})=\int p(\mathbf{c}^{*}|\mathbf{x}^{*},\mathbf{\{M}_{i}\}_{i=1}^{L})\\ q(\mathbf{\{M}_{i}\}_{i=1}^{L})d\mathbf{\{M}_{i}\}_{i=1}^{L}$ (1)
@@
$q(\mathbf{c}^{*}|\mathbf{x}^{*})=\mathcal{N}\big{(}\mathbf{c}^{*};\hat{\mathbf{c}}^{*}(\kappa(\mathbf{x}),\mathbf{z}_{1},\cdots,\mathbf{z}_{L}),\tau^{-1}\mathbf{I}_{R}\big{)}\\ \quad\text{Bern}(\mathbf{z}_{1}),\cdots,\text{Bern}(\mathbf{z}_{L})$ (2)
@@
$\hat{\mathbf{y}}^{*}=\sqrt{\frac{1}{r_{L}}}(\mathbf{M}_{L}\mathbf{z}_{L})\sigma\Big{(}\cdots\sqrt{\frac{1}{K_{1}}}\sqrt{\frac{1}{r_{1}}}(\mathbf{M}_{2}\mathbf{z}_{2}\sigma\big{(}(\\ \mathbf{M}_{1}\mathbf{z}_{1}\big{)}\kappa(\mathbf{x})^{*}+\mathbf{b}_{i})\cdots\Big{)}$ (3)
@@
where $\alpha$ and $\beta$ are the parameters of the Beta distribution. The parameter $\alpha$ represents the number of “successes” in a hypothetical experiment and $\beta$ represents the number of “failures” in the same experiment. Note that the choice of the Beta distribution is mainly due to conjugacy Fink (1997). A conjugate prior is a special type of prior distribution used in Bayesian inference, where it has a unique and convenient property: when combined with the likelihood function of the observed data, the resulting posterior distribution also belongs to the same family of distributions as the prior. This makes working with conjugate priors in Bayesian analysis, particularly advantageous.
@@
we can denote $\alpha_{D}=\sum_{d=1}^{D}x_{d}+\alpha$ and $\beta_{D}=D-\sum_{d=1}^{D}x_{d}+\beta$. Among the various kernel functions available, such as the radial basis kernel Patle and Chouhan (2013), our experiments yielded the best results with the squared kernel. This choice was guided by the squared kernel’s ease of differentiation and minimal computational overhead on the model.
@@ 4.1 Datasets
We utilize three open source and popular datasets in the medical field for our experiments, specifically the SOAP dataset Afzal et al. , the Medical Transcription dataset [^2] https://www.kaggle.com/datasets/tboyle10/medicaltranscriptions, and the ROND Clinical text classification dataset Liu et al. (2024). The SOAP dataset involves multi-class classification, encompassing four classes: subjective, objective, assessment, and plan. These classes are derived from the widely recognized medical protocol known as SOAP. For our experimentation, we employed their finalized dataset, which includes 152 clinical notes for training and 51 for testing. The Medical Transcription (MT) dataset is a multi-class classification dataset containing medical transcriptions for various medical specialities. This dataset exhibited an imbalance among the classes. For our experiments, we selected a subset of this dataset, which comprised the top four classes and included a total of 2,330 instances. Lastly, we used a subset of the publicly accessible Radiation Oncology NLP Database (ROND) known as clinical text binary classification of therapy type, comprising 100 cases.

### v2

# From Uncertainty to Trust: Kernel Dropout for AI-Powered Medical Predictions

@@ Abstract
AI-driven medical predictions with trustworthy confidence are essential for ensuring the responsible use of AI in healthcare applications. The growing capabilities of AI raise questions about their trustworthiness in healthcare, particularly due to opaque decision-making and limited data availability. This paper proposes a novel approach to address these challenges, introducing a Bayesian Monte Carlo Dropout model with kernel modelling. Our model is designed to enhance reliability on small medical datasets, a crucial barrier to the wider adoption of AI in healthcare. This model leverages existing language models for improved effectiveness and seamlessly integrates with current workflows. Extensive evaluations of public medical datasets showcase our model’s superior performance across diverse tasks. We demonstrate significant improvements in reliability, even with limited data, offering a promising step towards building trust in AI-driven medical predictions and unlocking its potential to improve patient care.
{IEEEkeywords}
Medical Text, Reliability, Language models, Kernel methods.
@@ 1 Introduction
Problem: “Alice, a medical professional at a busy healthcare firm, is tasked with managing numerous patient cases. To handle the growing workload, her firm implemented advanced AI solutions using powerful transformer models like Bio-BERT and Clinical-BERT. While these models promised efficiency by automating routine tasks, Alice found herself uncertain about their predictions, aware that even small errors could lead to significant consequences in the medical field. This uncertainty led her to manually review every prediction, increasing her workload instead of reducing it. Alice realized that what she needed was a system capable of assessing the model’s confidence in its predictions, enabling her to prioritize and focus on cases with lower confidence scores, thereby saving time and ensuring critical cases received the attention they deserved.”
@@
The current healthcare landscape is witnessing a remarkable transformation driven by machine learning [45, 10]. Automated methods are proliferating across diverse medical domains, empowering tasks like medical imaging analysis with the ability to detect anomalies [4], classify diseases like cancer [8], and even predict disease progression [61]. However, despite the undeniable power of these tools, responsible implementation necessitates careful consideration of ethical issues [55], data privacy [62], and the crucial role of human oversight. One persistent challenge lies in ensuring the consistently high reliability [23, 50] of automated predictions, a factor upon which the trust and adoption of these models by medical practitioners ultimately hinge [9, 24].
@@
Another challenge is developing computational methods that are reliable even on small datasets. In the medical domain, datasets are typically scarce, even those publicly available [51]. This scarcity stems from the limited release of critical private data, a trend driven by strong patient privacy concerns and regulations safeguarding that privacy in many countries [38]. Restricted access to data often forces training models on smaller datasets, making complete trust in their predictions more challenging. This underscores the critical importance of ensuring the reliability and dependability of these predictions. This paper addresses this challenge by presenting a novel computational model built on Bayesian Monte Carlo Dropout, with the specific objective of enhancing the reliability and trustworthiness of predictions made in the medical domain.
@@
Despite advancements, machine learning approaches in healthcare face challenges in building trust due to limitations in their transparency and reliability [12]. For example, deep learning, while powerful, often functions as a “black box”, making it difficult to understand its reasoning and assess its accuracy in individual cases. In contrast, deep Bayesian models [30, 1] provide a wide range of benefits [41, 30], for instance, in the medical predictions it could boast of several potential advantages [57, 21, 14]. A Bayesian model inherently provides uncertainty estimates with predictions, giving clinicians a clearer understanding of how confident the model is in its output. This allows for better-informed decision-making while acknowledging the limitations of the model. By averaging over multiple model configurations (due to dropout), the Bayesian model can be less susceptible to noise and biases present in training data, leading to more generalizable and reliable predictions. The Bayesian setting outputs represent probability distributions, allowing clinicians to understand the range of possible outcomes and the likelihood of each. This aids in explaining the model’s reasoning and building trust in its predictions. Analyzing dropout patterns can offer clues about which features influence the model’s predictions most, potentially revealing previously unknown relationships within the data. Bayesian modelling can be calibrated to ensure their predictions align with actual outcomes, further improving their trustworthiness.
@@
Studies have indicated that machine learning contributes to enhancing the efficiency of medical-related tasks [44]. However, there appears to be resistance among people when it comes to adopting medical artificial intelligence treatments [33]. Similarly, doctors also feel hesitant to fully embrace it due to the sensitive nature of the tasks and the opaque and black box nature of machine learning models  [^1] https://www.ntu.edu.sg/medicine/news-events/dean’s-blog/trust-can-we-trust-the-machine-can-we-trust-the-doctor [22]. From a larger perspective, ensuring reliable decisions is crucial, especially when someone’s life is on the line, allowing little room for mistakes. However, machine learning models face a challenge as they cannot measure the certainty of their decisions.
Our key contributions are the following:
  • We develop a novel computational model based on Bayesian deep learning that exploits a novel Bayesian Monte Carlo Dropout mechanism to model uncertainty.
@@
  • Our fully Bayesian model incorporates conjugate priors that allow us to naturally incorporate prior knowledge or beliefs about the parameters we are estimating. This can lead to more accurate and reliable posterior estimates, especially when data is limited.
  • We develop a kernel model to reliably model the features. Kernels can allow us to model the nature of the data and problem, allowing for better flexibility and adaptation.
  • We demonstrate that the novel model is reliable even on small datasets in binary and multi-class settings.
@@ 2 Related Work
The adoption of machine learning in the medical field is becoming increasingly widespread [7, 56], particularly with the development of transformer models tailored for medical applications [42, 28]. These models have notably improved the effectiveness and productivity of medical tasks [53, 39]. Especially there is rapid research going on in clinical judgment or diagnosis tasks in which the judgement is made based on clinical notes, medical transcriptions and the data based on the patient history [15, 19].
@@
In [63], the authors provided an overview of the current transformer models in the medical domain and their utilization in a range of medical applications, such as text mining, question answering, and classification. Similarly, [29] highlighted the challenges encountered in managing Electronic Health Records (EHRs) and how advanced natural language processing techniques are aiding in addressing these challenges. These techniques assist in organizing and generating the data, extracting valuable information, and facilitating classification and prediction tasks, thereby enhancing the effectiveness and efficiency of work processes. Considerable research is currently underway in this area, with [26] notably introducing a novel medical report generation model called TranSQ. This model generates a series of semantic features to align with plausible clinical concerns and constructs the report using sentence retrieval and selection techniques. Their model surpassed benchmark models in performance on the generation tasks. Additionally, [47] investigated the potential of the widely acclaimed Generative Pretrained Transformer 3 (GPT-3) and outlined considerations for its integration into clinical practices.
@@
Similar to other medical tasks, the classification task is also reaping the benefits of advanced machine learning techniques, aiming to improve healthcare and enhance patient care, given the importance of swiftly analyzing text for clinical decision support, research, and process optimization. [59] employed deep learning models to classify clinical records into five primary disease categories in traditional Chinese medicine. [3] performed multi-label classification of ICD-10 code problems by using transfer learning in conjunction with pre-trained BERT and BioBERT models. Meanwhile, [49] utilized a BERT model to identify diagnoses from unstructured clinical text. Similarly, [60] conducted multi-label classification of medical texts using advanced transformer models. Through their experiments, they concluded that domain-specific transformers yield superior results compared to standard transformers.
@@
There are certain constraints associated with employing these advanced transformer models for clinical text classification, as noted by [18]. One significant limitation is the black-box nature of these deep learning models, which poses challenges in trusting and implementing these models, particularly in healthcare, where decisions are matters of life and death [43]. These opaque models fail to provide a quantifiable measure of their output confidence. Research is underway to address this issue, with [17] employing dropout as a Bayesian approximation to represent uncertainty in deep learning models. [52] introduced two novel metrics, namely Verbalized Uncertainty and Probing Uncertainty, to measure the uncertainty of generated explanations by large language models.
@@ 3 Our Novel Model
This section dives into the technical details of our innovative model, which delivers a more reliable quantification of uncertainty in its predictions for multi-class classification. This uncertainty arises from the inherent lack of complete knowledge about the true relationship between input and output data. We address this uncertainty by differentiating our model from the standard Monte Carlo dropout model [17, 35] in two key ways.
@@
Our second innovation lies in integrating priors within the Monte Carlo dropout framework of [48]. Priors inject vital flexibility by allowing us to leverage our domain knowledge and beliefs about the problem at hand. This strategic infusion of information guides the model towards solutions that resonate with real-world expectations, effectively biasing it towards sensible outcomes. Additionally, incorporating priors empowers our Bayesian models to quantify uncertainty in their predictions. This explicit estimation of uncertainty is crucial for tasks where understanding the model’s limitations is paramount, such as in medical diagnosis or legal prediction. Finally, priors facilitate robust comparison and selection of different models based on their posterior probabilities. This data-driven approach empowers us to identify the model that most effectively aligns with both the observed data and our prior beliefs.
@@
Imagine a neural network constantly questioning its abilities. This is the essence of Monte Carlo dropout, a process where the network throws “curveballs” at itself during training. At each training step, the network randomly deactivates some neurons in its hidden layers, simulating “failures” and forcing others to take over. This introduces variability and prevents over-reliance on specific neurons. By “silencing” neurons, the network effectively asks, “Can I still work even without these?” This builds a more robust network, adaptable to new data and challenges. During prediction, the network performs multiple passes with different silenced neurons, mimicking multiple “runs” with potential errors. This captures the inherent uncertainty in real-world data and allows the network to express its level of confidence in each prediction. This connection between dropout and uncertainty estimation was not always clear. Earlier works like [17] focused on dropout as a regularization technique, but later studies (e.g., [25, 34]) revealed its deeper connection to Bayesian inference in Gaussian Processes. By leveraging dropout, we can not only improve a network’s performance but also gain valuable insights into its limitations and confidence in its predictions.
@@
$q(\mathbf{c}^{*}|\mathbf{x}^{*})=\int p(\mathbf{c}^{*}|\mathbf{x}^{*},\mathbf{\{M}_{i}\}_{i=1}^{L})\,q(\mathbf{\{M}_{i}\}_{i=1}^{L})\,d\mathbf{\{M}_{i}\}_{i=1}^{L}$ (1)
@@
$\begin{split}q(\mathbf{c}^{*}|\mathbf{x}^{*})=\mathcal{N}\big{(}\mathbf{c}^{*};\hat{\mathbf{c}}^{*}(\kappa(\mathbf{x}),\mathbf{z}_{1},\cdots,\mathbf{z}_{L}),\tau^{-1}\mathbf{I}_{R}\big{)}\\ \quad\text{Bern}(\mathbf{z}_{1}),\cdots,\text{Bern}(\mathbf{z}_{L})\end{split}$ (2)
@@
$\begin{split}\hat{\mathbf{y}}^{*}=\sqrt{\frac{1}{r_{L}}}(\mathbf{M}_{L}\mathbf{z}_{L})\sigma\Big{(}\cdots\sqrt{\frac{1}{K_{1}}}\sqrt{\frac{1}{r_{1}}}\\ (\mathbf{M}_{2}\mathbf{z}_{2}\sigma\big{(}(\mathbf{M}_{1}\mathbf{z}_{1}\big{)}\kappa(\mathbf{x})^{*}+\mathbf{b}_{i})\cdots\Big{)}\end{split}$ (3)
@@
where $\alpha$ and $\beta$ are the parameters of the Beta distribution. The parameter $\alpha$ represents the number of “successes” in a hypothetical experiment and $\beta$ represents the number of “failures” in the same experiment. Note that the choice of the Beta distribution is mainly due to conjugacy [16]. A conjugate prior is a special type of prior distribution used in Bayesian inference, where it has a unique and convenient property: when combined with the likelihood function of the observed data, the resulting posterior distribution also belongs to the same family of distributions as the prior. This makes working with conjugate priors in Bayesian analysis, particularly advantageous.
@@
we can denote $\alpha_{D}=\sum_{d=1}^{D}x_{d}+\alpha$ and $\beta_{D}=D-\sum_{d=1}^{D}x_{d}+\beta$. Among the various kernel functions available, such as the radial basis kernel [40], our experiments yielded the best results with the squared kernel. This choice was guided by the squared kernel’s ease of differentiation and minimal computational overhead on the model.
@@ 4.1 Datasets
We utilize three open source and popular datasets in the medical field for our experiments, specifically the SOAP dataset [2], the Medical Transcription dataset [^2] https://www.kaggle.com/datasets/tboyle10/medicaltranscriptions, and the ROND Clinical text classification dataset [32]. The SOAP dataset involves multi-class classification, encompassing four classes: subjective, objective, assessment, and plan. These classes are derived from the widely recognized medical protocol known as SOAP. For our experimentation, we employed their finalized dataset, which includes 152 clinical notes for training and 51 for testing. The Medical Transcription (MT) dataset is a multi-class classification dataset containing medical transcriptions for various medical specialities. This dataset exhibited an imbalance among the classes. For our experiments, we selected a subset of this dataset, which comprised the top four classes and included a total of 2,330 instances. Lastly, we used a subset of the publicly accessible Radiation Oncology NLP Database (ROND) known as clinical text binary classification of therapy type, comprising 100 cases.